OpenStudy (anonymous):

Help me solve tan 22.5 using the half angle theorem

OpenStudy (anonymous):

http://www.regentsprep.org/Regents/math/algtrig/ATT14/formulalesson.htm Let (1/2)(A) = 22.5 degrees Therefore, A = 45 degrees tan(0.5 * 45) = (1- cos45)/sin45 I know the value is positive because the angle is in the 1st quadrant: sin45 = cos45 = (sqrt2)/2 tan(0.5*45) = (1- (sqrt2)/2)/((sqrt2)/2)

OpenStudy (anonymous):

is it with 1+cos45 on the bottom?

OpenStudy (anonymous):

sin45. If you look at the website, the half angle formula for tan can be expressed in 3 different ways, which are all the same.

OpenStudy (anonymous):

so it would be \[\sqrt{1-\cos45\frac{ 1}{\cos 45 }}\]

OpenStudy (anonymous):

pretend that the thing is next to the 1/45

OpenStudy (anonymous):

so it would be \[1-\frac{ \sqrt{2} }{ 2 }\]

OpenStudy (anonymous):

\[1+\frac{ \sqrt{2} }{ 2 }\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

It would be: \[( 1-\sqrt{2}/2)/(\sqrt2/2)\]